### 26 Feb 2024 v1

# Neural Population Geometry and Optimal Coding of Tasks with Shared Latent Structure

Neural Population Geometry and Optimal Coding of Tasks with Shared Latent Structure
Albert J. Wakhloo Center for Theoretical Neuroscience, Columbia University Center for Computational Neuroscience, Flatiron Institute Will Slatton Center for Computational Neuroscience, Flatiron Institute Center for Neural Science, New York University SueYeon Chung Center for Computational Neuroscience, Flatiron Institute Center for Neural Science, New York University
1 Abstract
Humans and animals can recognize latent structures in their environment and apply this information to efficiently navigate the world. Several recent works argue that the brain supports these abilities by forming neural representations that encode such latent structures in flexible, generalizable ways. However, it remains unclear what aspects of neural population activity are contributing to these computational capabilities. Here, we develop an analytical theory linking the mesoscopic statistics of a neural population’s activity to generalization performance on a multi-task learning problem. To do this, we rely on a generative model in which different tasks depend on a common, unobserved latent structure and predictions are formed from a linear readout of a neural population’s activity. We show that three geometric measures of the population activity determine generalization performance in these settings. Using this theory, we find that experimentally observed factorized (or disentangled) representations naturally emerge as an optimal solution to the multi-task learning problem. We go on to show that when data is scarce, optimal codes compress less informative latent variables, and when data is abundant, optimal codes expand this information in the state space. We validate predictions from our theory using biological and artificial neural network data. Our results therefore tie neural population geometry to the multi-task learning problem and make normative predictions of the structure of population activity in these settings.
2 Introduction
Humans constantly solve different instances of similar problems. We brake at stop signs, stop at red lights, and slow down in crowded streets. We do these things effortlessly and efficiently learn to use new sensory cues to regulate our behavior. This is possible because we are able to recognize overt symbols like road signs, as well as more abstract visual cues like the crowdedness of a street. More generally, humans and animals learn to recognize latent variables in the environment and use them to guide their behavior across contexts and tasks.
Recent experimental findings have described coding strategies that may underlie this ability. In particular, several studies have described cases where independent variables in the environment are represented along distinct directions of variation in the neuronal activity space [1, 2, 3, 4, 5, 6]–e.g., in orthogonal subspaces of the firing rates of a collection of neurons [2, 7, 8]. These independent factors have ranged from the contacts of distinct mouse whiskers [2], to more abstract latent variables such as the values of different choices in a decision making task [3]. Neural representations in which distinct environmental variables are represented along independent or orthogonal directions of variation are referred to as factorized or disentangled. Factorized representations were recently shown to emerge in artificial networks trained on multiple tasks [7] and are thought to support generalization to new contexts as well as efficient learning of new tasks that depend on shared latent variables [4, 9].
Similarly, other studies have argued that the brain makes widespread use of cognitive maps to solve these problems [10]. These are coding strategies in which environmental variables are represented in the population code in a way that preserves task-relevant relations between them. This idea is supported by a range of findings, for example, studies in which structurally similar tuning profiles emerge in a neural population for distinct types of environmental variables [11, 12, 13, 14, 15]. These variables have ranged from an animal’s position [16], to the frequency of an auditory stimulus [11], to more abstract quantities like the amount of evidence accumulated in a decision making task [13]. Here and in the findings referenced above, neural populations represent latent structure in the environment in a way that supports a target behavior. However, defining measures for and understanding why certain neural activity patterns represent latent variables in task-efficient ways remains challenging [17].
A promising approach to tying neuronal activity patterns to such computational goals is to study the geometry of the neural responses [18]. Here, the overarching idea is to find which mesoscopic statistics of the population activity contribute to a macroscopic target computation or behavior. In this way, we can gain insight into neural computation without having to give a detailed account of microscopic single unit activity. For example, in the domain of invariant object recognition, recent works analytically tied coding efficiency [19, 20, 21] and few-shot generalization performance [22] to measurable statistics of the population activity. Recent works have also described a wide range of motor [23, 24, 25, 26], sensory [2, 6, 14, 27, 28], and decision making [3, 4, 12, 13] computations by studying the geometry of neural population responses.
In this study, we develop an analytical theory for learning multiple tasks depending on a common latent structure that directly ties the statistics of neural population responses to generalization performance. While several authors have used linear probes or heuristic geometric measures to analyze population activity in such tasks (e.g. [2, 4]), to the best of our knowledge, there is no cohesive theory that directly ties mesoscopic statistical features of neural activities to multi-task learning performance. To fill this gap, we analytically calculate the generalization error of an agent learning multiple tasks that depend on a common latent structure. To do this, we use a flexible model in which binary classification tasks are formed in an unobserved latent space, and an agent makes decisions by applying a linear readout to a set of neural population activities [7, 29, 30, 31, 32].
Using these results, we show how the error is completely determined by a small set of geometric statistics summarizing the relationship between the latent variables and neural responses. We leverage this theory to derive normative predictions describing the optimal geometry that a set of population responses should take. In this way, we show that disentangled representations naturally emerge as an optimal solution to the multi-task learning problem. Furthermore, we show how optimal codes compress less useful information when data is scarce and expand this information in the activity space when data is abundant. We go on to describe how this strategy can be traced back to measurable properties of the eigenspectrum of neuronal responses. Finally, we show that our theory accurately predicts the generalization error of readouts trained on artificial neural network (ANN) data as well as multi-unit recordings taken from the macaque ventral stream [33], and we demonstrate the use of our geometric measures as a data-analytic tool.
3 Results
3.1 Theory of Multi-task Learning
Figure 1: Schematic of the task and model setup using images from the d-sprites dataset as an example [34]. (a) Visual stimuli are generated from points in a latent space, and binary discrimination tasks are formed by linearly separating the latent space using a hyperplane with normal $T_{1}$. (b) Each stimulus elicits a neuronal activity pattern, visualized as points in an activity space. (c) A new binary discrimination task can be formed by separating the latent space using a different hyperplane with normal $T_{2}$. (d) As before, stimuli elicit neural activity patterns. (e-f) Schematic of the Hebbian learning rule we consider. Given a set of $p=6$ training stimuli, subsequent decisions are made using a linear readout of the neural activity patterns, in this case the firing rates of 3 neurons. When the number of positive (red) and negative (blue) labels is balanced, this linear readout corresponds to using a hyperplane whose normal points in the direction of the difference in the means of positive and negative examples .
Figure 2: Schematic of the geometric terms. We visualize different possible neuronal activity patterns elicited by the same set of stimuli. (a) A small slice of the latent space from which stimuli are generated. (b) Visualization of neural activity patterns with low (left) and high (right) total correlation, $c$. When the correlation is high, the relative distances between points in the latent space are preserved in the firing rate space. (c) Alignment term. When the alignment is low, different latent variables are represented along the same coding directions, and the noise distribution (grey ellipses) around a given point in the firing rate space falls along the coding directions. When the alignment is high, the noise distribution is orthogonal to these directions in the state space, and independent directions in the latent space are represented along approximately orthogonal coding directions in the neural state space. In other words, high alignment codes are factorized, with noise distributions that are orthogonal to signal directions. (d) Dimensionality term. In higher dimensional representations, the noise distribution occupies more directions in the state space. Here, we visualize this as a 3d noise distribution (right) as opposed to a 2d distribution (left).
We study the ability of a neural population to support downstream learning of tasks that depend on a common latent structure. Specifically, we consider a setting in which an agent learns binary classification tasks using a set of $p$ training stimuli. These stimuli are formed from a set of unobserved latent variables (Fig 1a). The task labels for a given binary classification task are formed by linearly separating the latent space into two pieces using a hyperplane with a normal vector, $T$. Note that while the separation is linear in the latent space, it may be highly non-linear in the stimulus space. Alongside the latent variables, we consider the activity patterns of $n$ neurons for each stimulus. We denote the vector of activity patterns for each stimuli as $x\in\mathbb{R}^{n}$. We quantify how well these neuronal activity patterns support downstream classification by calculating the generalization error of an agent that makes predictions using a linear readout of these activities, which is learned using a supervised Hebb rule (Methods; Fig. 1(e-f)). In this way, we connect the statistical properties of the neural code to the multi-task learning problem.
While the generalization error for arbitrary distributions of neuronal activities and latent variables is analytically intractable, we show that in many cases, the error only depends on a few key statistics. To do this, we draw from recent work in deep learning theory showing that the generalization error of linear readouts trained on complex distributions can in many cases be well approximated by studying simpler Gaussian models (Methods) [30, 35, 31, 36]. Thus, we analytically calculate the generalization error using a Gaussian model and show empirically that our theory accurately accurately predicts the generalization error of the linear readout rule when applied to non-Gaussian synthetic data (Sec. 3.3) as well as real neural responses from V4 and IT on a visual classification task (Sec. 3.4).
Using this simplified model, we prove a formula for the generalization error. As described in the Supplemental Material (SM), we show that under certain limiting conditions, the generalization error $E_{g}$ can be written as a strictly decreasing function of 3 geometric terms, which we schematize in Fig. 2. We now discuss each of these one by one:
  • Neural-Latent Correlation: $c.$ This term is a normalized sum of all covariances between neurons and latent variables. Geometrically, the total correlation measures how sensitive the population responses are to variations in the latent space.
  • Alignment: $a.$ This term measures the alignment between the coding directions for each latent variable and the total variation in the neural response. Because neural response variation reflects a combination of latent-driven variation (i.e., signal) and latent-independent noise, the alignment term favors encoding each latent variable along a direction that contains little signal from the other latent variables and which has low latent-independent noise. Thus, a high alignment $a$ reflects low noise in the coding directions and near orthogonal coding directions for each latent variable (i.e., a factorized code).
  • Neural Dimensionality: $\operatorname{PR}.$ The participation ratio of the neural responses measures the number of dimensions that the population activity spans. When all else is equal, higher dimensional responses are preferred, as neuronal noise is less correlated from trial to trial [22].
Finally, we note that the contribution to the error of neuronal dimensionality $\operatorname{PR}$ decays to 0 with the number of training samples (Methods). Similarly, we find that the correlation term’s contribution to the error decays with the number of samples. On the other hand, the error stemming from the alignment, $a,$ represents an irreducible error that does not depend on the number of training samples. Thus, we anticipate that as the number of training samples grow, it becomes increasingly important to maximize the alignment term, while in the few-shot regime, the correlation and participation ratio terms dominate the error.
Figure 3: Theory predicts empirical generalization error in Gaussian model with power law covariance spectra. (a) Schematic illustrating simulation setup. Gaussian latent variables $z$ are used to generate task labels, and predictions are formed using a linear transformation of the latent variables, $x=Az$. (b-d) Two typical units for the (b) latent variables, (c) random high dimensional projection, and (d) whitened transform for various values of the spectral decay exponent, $\alpha.$ (e) Eigenvalues of the latent covariance, $\Omega$, for different decay rates, $\alpha.$ (f-h) Multi-task generalization error as a function of training samples, $p$, for (f) the latent variables themselves, (g) the random projection, and (h) the whitened transform.
We begin by validating our theory numerically on data drawn from a Gaussian model. Here, we find that our theory yields an excellent agreement with numerical simulations for a wide range of values (Methods). In these simulations, we sample the latents $z_{\mu}$ and neural responses $x_{\mu}$ from a multivariate Gaussian. We set the covariance matrix of the latents to have a spectrum that decays as a power law. Since the proof of our main theorem assumes that the spectra of the covariances decays more slowly than a power law with rate 1, this allows us to parametrically study violations of our main theorem’s assumptions (Fig. 3a&e; SM). After sampling the latents, we form the neural responses either by taking a random high dimensional projection of the latent variables, or by applying a whitening transform to the latent variables (Fig. 3b-d; see Methods). We then calculate the empirical generalization error across a set of binary classification tasks which are formed by shattering the latent space as above.
As shown in Fig. 3f-h, our theory yields an excellent fit to numerical simulations. Importantly, the theoretical prediction holds all the way down to the few shot learning regime in which $p$ is small, despite the theory being derived in the limit of large $p$. Furthermore, we find that the theory predicts the empirical generalization error well for relatively small values of $n$ and $d$, as well as latent variables whose covariance has an eigenspectrum that decays faster than what our main theorem assumes ($\alpha<1$, where $\alpha$ is the rate of decay). These results suggest that our theory can be applied to a wide range of datasets and is informative of the few shot learning regime.
3.2 Optimal Representation of Latent Variables
Figure 4: Optimal representational geometry as a function of training samples and latent structure. (a) In our task setup, directions in the latent space that have little variance are, on average, less informative of the task labels (Methods). Optimal neuronal representations of these latent variables map independent directions in the latent space to independent directions in the neuronal space–i.e. they are factorized. The amount of variance corresponding to less informative directions in the latent space is small when data is scarce and is large when data is abundant. (b) Eigenvalues of the optimal neural covariance as a function of the number of samples vs. latent dimensionality, $p/d$. We show the eigenvalues of the latent variables’ covariance in black. Markers correspond to results obtained by optimizing our formula for the generalization error numerically, and solid lines correspond to our formula for the optimal code’s spectrum (Methods). As the number of samples increases, the spectrum of the optimal neural code becomes increasingly flat, indicating the expansion strategy described above. (d-f) Geometric terms of the optimal neural representation for various values of $p$ and using the same latent covariance as in (c). Surprisingly, the total correlation decreases with $p$.
Our theory provides an ideal framework to pose normative questions regarding multi-task learning. In our model, latent variables which have less variance contain, on average, less information about the task labels (Fig. 4a). This allows us to probe how task structure and the number of available training samples determine the optimal representation. In this section, we study what the best possible neural representational geometry is, given some latent variables and a fixed training dataset size.
We show analytically that the optimal representation disentangles latent variables into orthogonal subspaces. More precisely, we show that each orthogonal direction of variation in the latent space maps onto an orthogonal direction of variation in the neural state space. Thus, disentangled representations emerge as the optimal solution to the multi-task learning problem.
In addition to being disentangled, we find that the optimal code compresses less informative variables when data is scarce and expands these variables when data is abundant. Importantly, while each latent is represented in an orthogonal direction, the variance along each direction is not equal. As shown in Fig. 4a, when the training budget is small, less informative latent variables correspond to directions in the neuronal activity space that have small variance. As $p$ increases, the amount of variance dedicated to these less informative latent variables grows. This reflects a strategy in which the optimal code compresses less informative variables at small training set sizes $p$, while at large sample sizes, these variables are expanded in the state space (Methods). Speaking informally, the optimal code only starts paying attention to the less informative latent variables when there are enough samples to learn their relevance to a given task.
We trace these features of the optimal code back to the eigenvalues of the neuronal covariance matrix and our three geometric terms. As shown in Fig 4b, the eigenspectrum of the optimal code becomes increasingly flat as $p$ increases. This reflects the fact that more and more variance is being dedicated to the less informative directions in the latent space (i.e., directions with smaller variance). Turning to the geometric terms, we can see that tradeoff appears as a trend in which the participation ratio and alignment of the optimal code both increase with $p$, while the correlation decreases (Fig. 4c-e). The fact that high dimensional representations can be preferred for large $p$ is a surprising consequence of our calculation, as the neuronal dimensionality becomes less important to the error as $p$ grows (Methods). A key normative prediction from our theory is therefore that the spectrum of neural responses becomes increasingly high dimensional as an agent learns to perform tasks that depend on a complex latent structure.
3.3 Geometry of Multi-task Learning in MLPs
Figure 5: Theory predicts generalization error of the Hebb rule in trained and random MLPs. (a) Schematic of the experiment. Latent variables $z$ are randomly shattered to generate task labels. These latents are passed through a random MLP (light blue) and are then used as inputs to train a 3 hidden layer MLP (dark blue) on the multi-task binary classification problem using stochastic gradient descent. (b-c) After training, we sample a new set of latents and teacher vectors and calculate the generalization error of the Hebb rule on each layer of the (b) random and (c) trained network. Theoretical predictions closely track empirical errors, and the trained network achieves a lower error in intermediate layers. (d-f) Geometric terms across layers for the random (light blue line) and trained (dark blue line) networks. Linear layers are marked by circles and relu layers by squares. Interestingly, the error only slightly changes across linear and relu layers of the same model stage, in spite of sharp changes in the geometry. (g-i) Average change in the geometry between relu and linear layers. Here we show the mean and standard deviation of the difference between each geometric term before and after the relu nonlinearity is applied. The application of relu consistently causes increases in the dimension and alignment, as well as decreases in the correlation.
To test our theory on explicitly non-Gaussian data, we now apply our theory to both random and trained non-linear multi-layer perceptrons (MLPs). As shown in Fig. 5a, we first sample a set of zero-mean Gaussian latent variables $z$. We then feed these latent variables through a random MLP with increasing hidden layer sizes. After generating this set of stimuli, we sample $500$ random $T$ vectors, and use these to generate a set of task labels (Methods). These labels and data points from the random MLP are then used to train a downstream 4-hidden-layer MLP, which predicts the label for each task using a linear readout of its penultimate layer. This is reminiscent of the hidden manifold modeling framework studied in the machine learning community [30]. Finally, we sample a new set of latent variables along with a new set of tasks and calculate the generalization error of the Hebb rule when applied to the representations at different layers of the trained and untrained MLPs (Methods). This setup allows us to validate our theory on non-linear transformations of Gaussian latent variables.
As shown in Fig. 5b-c, we find good agreement between our theoretical predictions and the empirical generalization error. Examining the layer-by-layer generalization errors we find that the random MLP successfully tangles neural representations of the latents, from the perspective of the downstream Hebbian readout, as the generalization error increases through the random MLP. On the other hand, we find that the trained MLP successfully manages to learn the latent structure of the task, as demonstrated by the fact that the generalization error of the Hebb rule drops sharply through the layers of the network. Overall, these results provide compelling evidence that our analytical formula holds well beyond settings in which the neural responses are Gaussian variables.
Turning to our geometric decomposition of the error, we find several interesting trends across layers and non-linearities (Fig. 5d-f). In the random MLP, we see that the correlation decreases throughout the network, while the alignment term exhibits small fluctuations while gradually decreasing overall. Since we choose the layer sizes to be monotonically increasing through the network, we also find that the dimension of the responses stays more or less constant through the network.
In the trained network, we find that linear and relu layers orchestrate a trade off between dimensionality, alignment, and correlation that together only slightly perturb the generalization error. As shown in Fig 5g-i, the neural dimensionality and alignment spike each time the nonlinearity is applied, at the cost of the total correlation. Conversely, at each linear layer, the correlation sharply rises while the dimensionality and alignment fall. Importantly, this pattern is not present in the random network. Thus, the trained MLP learns to use the nonlinearity to increase the dimension of the representations while simultaneously squashing particularly harmful noisy directions of variability in the input and increasing how well factorized the code is. Intuitively, the linear layers learn to orient signal-unrelated features into parts of the state space that are zeroed out by the non-linearity [37]. While these are sharp changes in the geometry, the overall generalization error across layers exhibits only minor fluctuations between most linear and relu layers, highlighting the limitations of a generalization error-based approach to analyzing network activity.
We now study the dynamics of these geometric terms through training. To do this, we track the layer-wise generalization error and geometry through a single epoch of training (Methods). As shown in Fig. 6, we find that the generalization error in this context decays through training across layers. In the linear layers, we find a small increase in the total correlation and dimension as well as a steady rise in the alignment. On the other hand, in relu layers, improved generalization is driven by an increase in the dimension and alignment across both early and late layers. In these layers, the total correlation decreases throughout training. We find similar trends for MLPs with a $\tanh$ non-linearity (SM). These geometric trends mirror the effects of adding more samples on the optimal representation: As more samples become available, the alignment and dimension grow while the correlation shrinks. In other words, the dynamics of MLP representations in non-linear layers qualitatively mimic the behavior of the optimal neural code described by our theory.
Figure 6: Evolution of generalization error and representational geometry through training. (a) Generalization error of the Hebb rule applied to four different layers on a previously unseen set of tasks. We examine layers from the early stages of the MLP (black) as well as late stages (green) for both linear (solid line) and relu (dashed line) layers. We can see that the error decreases across all layers with the number of training samples, $p$. (b) Dimension term. While the dimension of the neural activations increases for all layers through training, the differences are most pronounced in the relu layers. (c) Correlation term. This term decreases through training for relu layers and slightly increases for linear layers. (d) Alignment term. This uniformly increases through training across the network. These findings mirror the effect of increasing the number of training samples on the optimal neural code.
3.4 Predicting Readout Performance of Macaque Visual Representations
Having considered non-linear artificial neural networks, we now apply our theory to biological neural data. To do this, we draw from pre-existing multi-unit recordings from macaque V4 and IT taken while 2 monkeys viewed visual stimuli (Methods; Fig. 7(a-c)) [33]. The stimuli used in these experiments included images of 64 objects taken from 8 categories and were generated by modifying $d=6$ continuous latent variables, that included the size, position, and angle of the object. This allows us to form binary classification task labels by shattering this 6-dimensional latent space on subsets of the data corresponding to individual object categories–e.g., a given task could involve separating images of chairs on the left side of the screen from those on the right. We test our theory by calculating the generalization error of the Hebbian readout rule applied to the V4 and IT neural responses.
We find that our formula accurately predicts the generalization error of the monkey neural responses across both recording sites. Furthermore, as shown in Fig. 7d, both V4 and IT achieve better generalization error than the raw pixels, and IT achieves a better generalization error than V4. [38]. Applying our geometric decomposition of the generalization error to these data, we find that the dimensionality is lower in V4 than in either IT or the pixel data [22] (Fig. 7e). Furthermore, we find that the alignment is far lower in V4 than in IT, suggesting that factorization of these latent variables increases along the ventral stream, in line with previous results [39]. However, we find that the alignment is higher in the pixels than in either V4 or IT. This finding suggests that the ventral stream representations considered here do not factorize information regarding object position well, as a poor alignment indicates that neuronal noise and the coding directions of distinct latents are not orthogonal to one another (Fig. 7e-g). Nevertheless, we do find that the total correlation is much higher in V4 and IT than in the pixel space. This suggests that while the representations considered here are not well factorized, they do contain correlations with the latent variables that can be leveraged to achieve a generalization error that is better than classification on the pixels. Taken together, these results provide strong support for our theory and demonstrate the applicability of our metrics as a tool for tying the geometry of neuronal population responses to the computational objective of multi-task learning.
Figure 7: Theory predicts multi-task error in macaque V4 and IT data. (a-b) Example stimuli and tasks. Visual stimuli included 64 objects grouped into 8 categories and were generated by modifying $d=6$ continuous latent variables that included object size, position, and angle. We form binary classification tasks on subsets of the data coming from the same category–e.g. all images from the ”Tables” category. (c) Probe placement. Figure adapted with permission from [33]. (d) Generalization error across pixels and brain regions calculated empirically (markers) and using our formula for the generalization error (solid line). (e-g) Geometric terms across pixels and neural responses. We can see that V4 representations are lower dimensional than either the pixels or IT, and that both IT and V4 suffer from a lower alignment than the pixels. However, both regions present with a much higher total correlation with the latent variables, enabling improved generalization performance.
4 Discussion
In this work, we analyzed a model of learning multiple tasks that share a common latent structure, connecting neural population geometry to the multi-task learning problem. In this model, binary classification tasks are generated by shattering a latent space uniformly. By varying the spectrum of the latent covariance, we were able to model situations in which certain latent variables are, on average, more useful for downstream tasks than others. We then calculated the generalization error of a linear readout applied to a set of neural activations responding to stimuli that were generated from these latent variables. In this way, we evaluated how well a neural representation is able to extract information about latent variables from the stimuli and use this for arbitrary downstream tasks.
To connect the mesoscopic statistics of neural activation patterns to generalization performance, we decomposed our formula for the generalization error into three geometric terms. These terms summarize the relationship between population activity and the latent variables and completely determine generalization performance in our task. First is the dimensionality of the entire population activity, as measured by the participation ratio. Second, we introduced the total correlation between neuronal activity and the latent variables, representing the overall sensitivity of neuronal responses to changes in the latent space. Finally, we defined the alignment, which summarizes the degree to which distinct latent variables and latent-unrelated variability are coded along orthogonal directions in the neural state space. As such, this term measures the degree to which a representation is factorized and manages to keep noise variability out of the coding directions. Together, these three terms completely characterize the generalization performance of the linear readout applied to the neural population responses.
To investigate the role that latent structure and data scarcity play in shaping population codes, we next analytically calculated which geometries minimize the generalization error, given a fixed latent structure and number of available training samples. This falls in line with studies on efficient coding in which predictions of sensory physiology are derived from statistics of the natural environment [40, 41, 42]. In this way, we found that disentangled (or factorized) representations naturally emerge as an optimal solution to the multi-task learning problem. Specifically, we found that the eigenvectors of the neuronal and latent covariances match the singular vectors of the cross-covariance (Methods). This is reminiscent of work done on the successor representation and grid cells [43, 44, 45, 46, 47], and it would be interesting to see if there are additional constraints under which grid cells or other coding properties emerge as an optimal code in our model [48].
We then described how the optimal neuronal code shifts from being low to high dimensional as the number of available samples increases. This finding reflects a strategy of compressing less useful information when data is scarce and expanding it when data is abundant. It is possible that a similar relation holds for low vs. high noise learning scenarios [49, 50, 51, 52]. It is interesting to compare these findings to previous work [22] reporting that higher dimensional representations of individual object classes are preferred for invariant object recognition in the few shot learning setting. Although the contribution to the error of the dimensionality scales as 1/ p in our formula just as in this work, the additional contribution to the error of the alignment term, leads to optimal representations with lower dimensionality in the few shot regime. These results highlight the fact that different geometric terms may compete with each other in ways that cannot be directly read off from generalization error equations when there are additional constraints imposed on the system. Finally, we note that theories of least squares estimation suggest that similar phenomena occur for more complex readouts [53, 54, 55, 56]. We suspect that the same spectral expansion trend we describe holds for such readouts, though perhaps the optimal code can remain smooth in such cases [27]. It would be interesting to investigate these questions either analytically or numerically in future work.
To test our theory, we next considered a setting in which a set of Gaussian latents are fed through a random network, and task labels are generated as above in the latent space. A four hidden layer MLP was then trained on the outputs of the random network on these tasks [30]. We then evaluated each layer of the trained network on previously unseen tasks using the scheme described above. Here, we found that intermediate layers of the trained network successfully learned to represent the latent variables in a way that could be used for downstream tasks, in line with previous work [7]. Furthermore, we found a strong correspondence between our theoretical predictions and the empirical generalization errors. Applying our geometric decomposition of the error, we found that linear and relu layers obtained similar generalization errors but with sharply different underlying geometries. Specifically, relu layers yielded a large boost to dimensionality and alignment at the cost of the total correlation. This suggests a strategy in which linear layers attempt to place neuronal noise in regions of the state space that are then squashed by the non-linearity [37]. Tracking the evolution of these terms through training, we found that the correlation in relu layers decreases through training, while the alignment and dimension increase. This mirrors the behavior of an optimal code as more samples are presented as well as the behavior of deep linear networks trained on certain regression tasks [57, 58]. It would be interesting to apply our geometric decomposition to analyze disentanglement in more complex artificial networks.
Finally, we applied our theory to electrophysiological recordings from macaque V4 and IT [33]. Specifically, we used our theory to predict the generalization error of a linear readout of neural responses trained on a series of binary classification tasks. We formed task labels using latent variables describing object position, size, and orientation. As before, our theory showed excellent agreement with empirical generalization errors of linear readouts applied to the neural responses. We found that both regions showed better readout performance than the pixels themselves [38]. While we found a relatively low alignment in these neural responses, we did find that alignment improved from V4 to IT. These findings suggest that, while these latent variables are not overall well factorized in the ventral stream, the degree of factorization improves from V4 to IT, partially in line with previous results [39]. It would be very interesting to repeat this experiment with visual regions specialized for localization. Overall, these results demonstrate the applicability of our theory as a data-analytic tool.
While we primarily focused on neural representations, it would be very interesting to derive behavioral predictions from this theory. Specifically, one could decompose the generalization error formula into error rates for easy vs. hard examples. In this way, we could leverage our theory to predict learning curve dynamics across stimuli of varying levels of difficulty. These predictions could then be tested in perceptual discrimination tasks involving continuous modulation of an unobserved latent space. We hope to pursue this line of research in subsequent work.
5 Methods
5.1 Model of Multi-task Learning
We model a setting in which an agent learns a set of binary classification tasks by performing a linear readout on a set of neural population activities. Formally, we assume that each stimulus is associated with a $d$-dimensional latent variable $z_{\mu}\in\mathbb{R}^{d}$, with $1\leq\mu\leq p$ denoting the sample index. The labels for a specific task are formed by shattering the latent space using a hyperplane with a normal vector, $T$. Thus, we have that the binary classification task labels, $y_{\mu},$ satisfy the relation, $y_{\mu}=\mathrm{sign}(T\cdot z_{\mu})$, so that each $T$ vector defines a specific classification task. Associated with each latent, we also consider $n$-dimensional neural activity patterns, $x_{\mu}\in\mathbb{R}^{n}$. From these data, the agent then forms predictions of new data points using a supervised Hebbian readout rule [59, 32]. More precisely, given a new stimulus associated with latent variables $z_{+}$ and firing rates $x_{+}$, the agent forms a prediction $\hat{y}_{+}$ using the rule
$\hat{y}_{+}=\mathrm{sign}(w\cdot x_{+}),\quad w=\frac{1}{p}\sum_{\mu}y_{\mu}x_{\mu}.$ (1)
When the labels are balanced, this corresponds to using the difference in the mean activities for positively and negatively labeled examples (Fig. 1f). We evaluate how well the neural code can support downstream classification by calculating the generalization error of these predictions across different tasks (i.e., different $T$ vectors).
While we validate our theory on a wide range of data, we obtain our analytical results using a simplified Gaussian model. This is inspired by work on the Gaussian Equivalence Principle (GEP) in the deep learning theory community [35, 30, 31]. Formally, we assume that each pair of neural responses and latent variables, $(x,z)$ are jointly zero-mean Gaussians with covariance matrices:
$\mathbb{E}[xx^{\top}]=\Psi,\quad\mathbb{E}[xz^{\top}]=\Phi,\quad\mathbb{E}[zz^{\top}]=\Omega.$ (2)
We can see that $\Psi$ describes the neuron-neuron covariances, $\Phi$ contains the covariances between single-unit responses and variations in the latent variables, and $\Omega$ describes covariances between latent variables in the dataset. Note that since the $T$ vectors are chosen randomly from a Gaussian distribution, the latent covariance determines which directions in the latent space are, on average, most informative of the task labels. Directions in the latent space which have a significant amount of variance are typically more informative of the task labels in this setup. This model corresponds to a variant of the popular student-teacher model [29, 31]. Insofar as the GEP holds, the neuronal and latent (cross-)covariances fully specify the generalization error of the linear readout.
5.2 Geometric Decomposition of Generalization Error Across Tasks
We prove a formula for the generalization error of the linear readout, averaged across different binary classification tasks (i.e., different $T$ vectors). To do so, we consider the limiting case in which the number of neurons, latent dimensions, and training samples are all large and of comparable size. Furthermore, we assume that the covariance matrices given above satisfy certain spectral properties–essentially we assume that their spectra decay at least as slowly as a power law with rate $\alpha<1$ (SM). Under these assumptions, we show that the average generalization error, $E_{g}$, is a decreasing function of the three geometric terms. Formally, we have
$E_{g}=\frac{1}{\pi}\tan^{-1}\bigg{(}\sqrt{\frac{\pi}{2pc^{2}\mathrm{PR}(\Psi)}+\frac{1}{ca}-1}\bigg{)},$ (3)
where we have introduced the total neural-latent correlation $c$, the alignment $a$, and the dimensionality of the population responses as measured by the participation ratio, $\mathrm{PR}(\Psi)$. These three terms correspond to statistics of the covariance matrices specified above:
$c$ $=\frac{\mathrm{Tr}(\Phi\Phi^{T})}{\mathrm{Tr}(\Psi)\mathrm{Tr}(\Omega)}$ (4)
$a$ $=\frac{\mathrm{Tr}(\Phi\Phi^{T})\mathrm{Tr}(\Psi)}{\mathrm{Tr}(\Phi^{T}\Psi\Phi)}$ (5)
$\operatorname{PR}(\Psi)$ $=\frac{\mathrm{Tr}(\Psi)^{2}}{\mathrm{Tr}(\Psi^{2})}$ (6)
Since the function $F(w)=\frac{1}{\pi}\tan^{-1}\big{(}\sqrt{x-1}\big{)}$ is strictly increasing, we can see that each of these three geometric terms should be made as large as possible. Note that in writing Eq. (3), we have separated a term that depends on the number of training samples, $1/[c^{2}\operatorname{PR}(\Psi)]$ from a term which is independent of the sample size, $1/(ca)$. Thus we can see that the correlation and dimensionality terms become less important as $p$ grows.
5.3 Gaussian Simulations
To test our theory on data points which violate the assumptions of our theory, we began by sampling from a finite Gaussian model. Specifically, we drew latent variables, $z_{\mu}$, from a multivariate Gaussian distribution whose covariance matrix had eigenvalues that decayed as a power law with rate $\alpha$. Specifically, we set: $\omega_{i}=5i^{-\alpha}$, where $\alpha$ was the power law of the spectrum, and $\omega_{i}$ is the $i$th eigenvalue of the latent covariance. The neural responses were then given by the formula $x_{\mu}=Az_{\mu}$ for a random, $n\times d$ Gaussian matrix $A$ with i.i.d. elements, or by applying a whitening transform $x_{\mu}=\Omega^{-1/2}z_{\mu}$. We set $n=80$ and $d=40$ for these simulations. For a fixed training set, we sampled $N_{task}=300$ task $T$ vectors and calculated the generalization error across all tasks using a set of new latent variables. Finally, we averaged over this entire procedure 30 times to generate the markers in Fig 3.
5.4 Optimal Codes
We derive which neuronal codes achieve the lowest generalization error, given a fixed number of samples to train on and a fixed latent structure. Specifically, we calculate which neuron-neuron and neuron-latent covariance matrices, $\Psi$, $\Phi$ achieve the lowest multi-task generalization error, given fixed a fixed latent covariance $\Omega$ and training set size $p$ (SM). We do this by optimizing Eq (3) with respect to $\Psi$ and $\Phi$ subject to the constraint that the entire covariance matrix between neurons and latents be positive semi-definite, a necessary condition for the code to be realizable. Using this approach, we find that the left and right singular vectors of $\Phi$ are the eigenvectors of $\Psi$ and $\Omega$, respectively for the optimal code. This shows that independent directions in the latent space map directly onto independent directions in the firing rate space. Furthermore, we obtain the following simple formula for the eigenvalues of the optimal neural code. Denoting the eigenvalues of $\Psi$ and $\Omega$ as $\psi_{i}$ and $\omega_{i}$ respectively, we have up to a permutation symmetry:
$\psi_{i}=C\frac{\omega_{i}}{2p\omega_{i}+\sum_{k}\omega_{k}},$ (7)
where $C$ is an arbitrary constant. As $p$ grows, we can see that the spectrum becomes flatter, reflecting the expansion strategy, while as $p$ shrinks, the spectrum decays faster and faster, reflecting the fact that less informative directions are compressed in the state space (Fig. 4c).
To validate our calculation, we numerically calculated the optimal code by optimizing on the space of positive semi-definite matrices. Since the full covariance matrix is positive semi-definite, there must exist matrices $X_{1}$ and $X_{2}$ such that
$L=\begin{pmatrix}\Omega^{1/2}&0\\ X_{1}&X_{2}\end{pmatrix},\quad LL^{T}=\begin{pmatrix}\Omega&\Phi^{T}\\ \Phi&\Psi\end{pmatrix}$ (12)
The space of possible $X$ matrices is unconstrained, so we simply optimize Eq (3) with respect to the $X_{i}$ matrices and calculate $\Psi$ and $\Phi$ after the fact.
5.5 MLP Experiments
We used random and trained MLPs to test several predictions from our theory using explicitly non-Gaussian artificial neural response data. To generate these data, we first sampled a set of $d=40$ dimensional latent variables from a multivariate Gaussian distribution with eigenvalues $\omega_{k}=k^{-0.2}$, though we also show similar results for a flat spectrum (SM). For these experiments, we sampled $5\cdot 10^{5}$ latent variables independently from this distribution as a training dataset. Using these latent variables we generated a set of $N_{t}=500$ tasks by randomly shattering the latent space. Latent variables were then passed through a $3$ layer perceptron with randomly initialized weights. We chose the size of each intermediate layer to be twice the size of the previous one to ensure the dimension of the representation did not decrease. Each layer was composed of a linear transform, followed by a relu non-linearity (see SM for results using a tanh non-linearity). After sampling the task labels and passing the latents through the random MLP, we trained a $4$ hidden-layer MLP to predict task labels from the random MLP responses. The trained network was similarly composed of linear-batchnorm-relu blocks, and we used the adam optimizer to train the network [60] through a single epoch. This setup corresponds to a multi-task version of the hidden manifold modeling framework studied in the machine learning community [30].
We evaluated the generalization error through layers and training using a new set of latents and tasks. Specifically, we sampled $10^{3}$ latent variables from the same distribution as well as a new set of $300$ randomly selected tasks. To calculate the theoretical generalization error , we then used the network representations of these new latent variables to calculate the geometric metrics and evaluate the theoretical generalization errors using Eq. (3) with $p=300$. To evaluate the empirical generalization error, we randomly split the new set of latents into a set of 300 training points and 700 test points. This process was repeated to calculate the average test error of the Hebb rule across each layer and time point in training. Finally, we averaged over train/test splits to generate the markers shown in Fig. 5.
5.6 Macaque Analyses
We drew from a publicly available dataset containing multi-unit recordings from V4 and IT taken from 2 monkeys [33]. These recordings were taken as the monkeys viewed visual stimuli as described in [33] and contained 88 V4 and 168 IT neural sites. The stimuli for this task were drawn from a generative model in which a total of 64 objects coming from 8 categories were displayed against varying backgrounds. These images were generated by varying $d=6$ continuous latent variables that controlled the object size, angle, and position. Latent variables were drawn iid from a mean zero uniform distribution.
We tested the linear decodability of these latent variables from the neural firing rates using the scheme described in Sec. 5.1. For the raw pixels, we first carried out a Gaussian random projection onto a $5,000$ dimensional space before applying this scheme. Specifically, for each task condition and for each of the 8 object category types, we formed task labels for $N_{t}=300$ tasks by randomly shattering the latent space. For each condition and category type, there were $320$ examples per condition and object category type. We z-scored the latent variables to ensure that there was no especially informative latent variable. We then formed predictions using the supervised Hebbian readout described in Sec. 5.1 after mean centering the neuronal firing rates. This procedure was repeated over 15 different train/test data splits. To generate the results in Fig. 7, we then averaged the generalization error across the $N_{t}$ classification tasks, data splits, conditions, and image categories. Error bars denote the standard error of the mean for the estimated generalization error across the $c=8$ category types. In the SM, we also show the geometry and generalization errors for each of the categories individually, as well as the error obtained by pooling all categories together.
Acknowledgements
We thank Thomas Yerxa for comments on this manuscript. This work was funded by the Center for Computational Neuroscience at the Flatiron Institute of the Simons Foundation and the Klingenstein-Simons Award to S.C. All experiments were performed on the Flatiron Institute high-performance computing cluster.

### 11 Apr 2024 v2

@@
 Humans and animals can recognize latent structures in their environment and apply this information to efficiently navigate the world.
-Several recent works argue that the brain supports these abilities by forming neural representations that encode such latent structures in flexible, generalizable ways.
-However, it remains unclear what aspects of neural population activity are contributing to these computational capabilities.
-Here, we develop an analytical theory linking the mesoscopic statistics of a neural population’s activity to generalization performance on a multi-task learning problem.
-To do this, we rely on a generative model in which different tasks depend on a common, unobserved latent structure and predictions are formed from a linear readout of a neural population’s activity.
-We show that three geometric measures of the population activity determine generalization performance in these settings.
-Using this theory, we find that experimentally observed factorized (or disentangled) representations naturally emerge as an optimal solution to the multi-task learning problem.
-We go on to show that when data is scarce, optimal codes compress less informative latent variables, and when data is abundant, optimal codes expand this information in the state space.
-We validate predictions from our theory using biological and artificial neural network data.
-Our results therefore tie neural population geometry to the multi-task learning problem and make normative predictions of the structure of population activity in these settings.
+However, it remains unclear what aspects of neural activity contribute to these computational capabilities.
+Here, we develop an analytical theory linking the geometry of a neural population’s activity to the generalization performance of a linear readout on a set of tasks that depend on a common latent structure.
+We show that four geometric measures of the activity determine performance across tasks.
+Using this theory, we find that experimentally observed disentangled representations naturally emerge as an optimal solution to the multi-task learning problem.
+When data is scarce, these optimal neural codes compress less informative latent variables, and when data is abundant, they expand these variables in the state space.
+We validate our theory using macaque ventral stream recordings.
+Our results therefore tie population geometry to multi-task learning.
 2 Introduction
@@
 Recent experimental findings have described coding strategies that may underlie this ability.
-In particular, several studies have described cases where independent variables in the environment are represented along distinct directions of variation in the neuronal activity space [1, 2, 3, 4, 5, 6]–e.g., in orthogonal subspaces of the firing rates of a collection of neurons [2, 7, 8].
+In particular, several studies have described cases where independent variables in the environment are represented along distinct directions of variation in the neuronal activity space [1, 2, 3, 4, 5, 6, 7]–e.g., in orthogonal subspaces of the firing rates of a collection of neurons [2, 7, 8, 9, 10].
 These independent factors have ranged from the contacts of distinct mouse whiskers [2], to more abstract latent variables such as the values of different choices in a decision making task [3].
@@
 These are coding strategies in which environmental variables are represented in the population code in a way that preserves task-relevant relations between them.
-This idea is supported by a range of findings, for example, studies in which structurally similar tuning profiles emerge in a neural population for distinct types of environmental variables [11, 12, 13, 14, 15].
+This idea is supported by a range of findings, for example, studies in which structurally similar tuning profiles emerge in a neural population for distinct types of environmental variables [13, 14, 15, 16, 17, 18].
 These variables have ranged from an animal’s position [19], to the frequency of an auditory stimulus [13], to more abstract quantities like the amount of evidence accumulated in a decision making task [15].
@@
 However, defining measures for and understanding why certain neural activity patterns represent latent variables in task-efficient ways remains challenging [20].
-A promising approach to tying neuronal activity patterns to such computational goals is to study the geometry of the neural responses [18].
+A promising approach to tying neural activity patterns to such computational goals is to study the geometry of the neural responses [21].
 Here, the overarching idea is to find which mesoscopic statistics of the population activity contribute to a macroscopic target computation or behavior.
 In this way, we can gain insight into neural computation without having to give a detailed account of microscopic single unit activity.
-For example, in the domain of invariant object recognition, recent works analytically tied coding efficiency [19, 20, 21] and few-shot generalization performance [22] to measurable statistics of the population activity.
-Recent works have also described a wide range of motor [23, 24, 25, 26], sensory [2, 6, 14, 27, 28], and decision making [3, 4, 12, 13] computations by studying the geometry of neural population responses.
-In this study, we develop an analytical theory for learning multiple tasks depending on a common latent structure that directly ties the statistics of neural population responses to generalization performance.
+For example, in the domain of invariant object recognition, recent works analytically tied coding efficiency [22, 23, 24, 25] and few-shot generalization performance [26] to measurable statistics of the population activity.
+Recent works have also studied a wide range of motor [27, 28, 29, 30, 31], sensory [2, 6, 16, 32, 33, 34], and decision making [3, 4, 14, 15] computations and behaviors by analyzing the geometry of neural population responses.
+In this study, we develop an analytical theory for learning binary decision making tasks depending on a common latent structure that directly ties the statistics of neural population responses to generalization performance.
 While several authors have used linear probes or heuristic geometric measures to analyze population activity in such tasks (e.g.
-[2, 4]), to the best of our knowledge, there is no cohesive theory that directly ties mesoscopic statistical features of neural activities to multi-task learning performance.
-To fill this gap, we analytically calculate the generalization error of an agent learning multiple tasks that depend on a common latent structure.
+[2, 4]), to the best of our knowledge, there is no cohesive theory that directly ties mesoscopic statistical features of neural activities to task performance in these settings.
+To fill this gap, we analytically calculate how neural population geometry shapes the generalization error of an agent learning multiple tasks that depend on a common latent structure.
 To do this, we use a flexible model in which binary classification tasks are formed in an unobserved latent space, and an agent makes decisions by applying a linear readout to a set of neural population activities [8, 35, 36, 37, 38].
@@
 Furthermore, we show how optimal codes compress less useful information when data is scarce and expand this information in the activity space when data is abundant.
-We go on to describe how this strategy can be traced back to measurable properties of the eigenspectrum of neuronal responses.
+We go on to describe how this strategy would be reflected in measurable properties of the eigenspectrum of neuronal responses.
 Finally, we show that our theory accurately predicts the generalization error of readouts trained on artificial neural network (ANN) data as well as multi-unit recordings taken from the macaque ventral stream [39], and we demonstrate the use of our geometric measures as a data-analytic tool.
@@
 3.1 Theory of multi-task learning
+We study the ability of a neural population to support downstream learning of tasks that depend on a common latent structure.
+Specifically, we consider a setting in which an agent learns binary classification tasks using a set of $p$ training stimuli.
+These stimuli are formed from a set of unobserved $d-$dimensional latent variables, $z\in\mathbb{R}^{d}$ (Fig 1a).
+The task labels for a given binary classification task are formed by linearly separating the latent space into two pieces using a hyperplane with a normal vector, $T$.
+Note that while the separation is linear in the latent space, it may be highly non-linear in the stimulus space.
+Alongside the latent variables, we consider the activity patterns of $n$ neurons for each stimulus.
+We denote the vector of activity patterns for each stimuli as $x\in\mathbb{R}^{n}$.
+We quantify how well these neuronal activity patterns support downstream classification by calculating the generalization error of an agent that makes predictions using a linear readout of these activities, which is learned using a supervised Hebb rule (Methods; Fig.
+1(e-f)) [38].
+As described in the Supplemental Material (SM), we calculate the generalization error of this readout both for a single, fixed task as well as the average generalization error across all possible tasks.
+In this way, we connect the statistical properties of the neural code to the multi-task learning problem.
+While the generalization error for arbitrary distributions of neuronal activities and latent variables is analytically intractable, we show that in many cases, the error only depends on a few key statistics.
+To do this, we draw from recent work in deep learning theory showing that the generalization error of linear readouts trained on complex distributions can in many cases be well approximated by studying simpler Gaussian models (Methods) [36, 37, 40, 41].
+Thus, we analytically calculate the generalization error using a Gaussian model and show empirically that our theory accurately predicts the generalization error of the linear readout rule when applied to the activations of non-linear multi-layer perceptrons (MLPs; Sec.
+3.3) as well as real neural responses from V4 and IT on visual classification tasks (Sec.
+3.4).
 Figure 1: Schematic of the task and model setup using images from the d-sprites dataset as an example [42].
@@
 (e-f) Schematic of the Hebbian learning rule we consider.
-Given a set of $p=6$ training stimuli, subsequent decisions are made using a linear readout of the neural activity patterns, in this case the firing rates of 3 neurons.
+Given a set of $p=12$ training stimuli, subsequent decisions are made using a linear readout of the neural activity patterns, in this case the firing rates of 3 neurons.
 When the number of positive (red) and negative (blue) labels is balanced, this linear readout corresponds to using a hyperplane whose normal points in the direction of the difference in the means of positive and negative examples .
@@
 (a) A small slice of the latent space from which stimuli are generated.
-(b) Visualization of neural activity patterns with low (left) and high (right) total correlation, $c$.
-When the correlation is high, the relative distances between points in the latent space are preserved in the firing rate space.
-(c) Alignment term.
-When the alignment is low, different latent variables are represented along the same coding directions, and the noise distribution (grey ellipses) around a given point in the firing rate space falls along the coding directions.
-When the alignment is high, the noise distribution is orthogonal to these directions in the state space, and independent directions in the latent space are represented along approximately orthogonal coding directions in the neural state space.
-In other words, high alignment codes are factorized, with noise distributions that are orthogonal to signal directions.
-(d) Dimensionality term.
-In higher dimensional representations, the noise distribution occupies more directions in the state space.
-Here, we visualize this as a 3d noise distribution (right) as opposed to a 2d distribution (left).
-We study the ability of a neural population to support downstream learning of tasks that depend on a common latent structure.
-Specifically, we consider a setting in which an agent learns binary classification tasks using a set of $p$ training stimuli.
-These stimuli are formed from a set of unobserved latent variables (Fig 1a).
-The task labels for a given binary classification task are formed by linearly separating the latent space into two pieces using a hyperplane with a normal vector, $T$.
-Note that while the separation is linear in the latent space, it may be highly non-linear in the stimulus space.
-Alongside the latent variables, we consider the activity patterns of $n$ neurons for each stimulus.
-We denote the vector of activity patterns for each stimuli as $x\in\mathbb{R}^{n}$.
-We quantify how well these neuronal activity patterns support downstream classification by calculating the generalization error of an agent that makes predictions using a linear readout of these activities, which is learned using a supervised Hebb rule (Methods; Fig.
-1(e-f)).
-In this way, we connect the statistical properties of the neural code to the multi-task learning problem.
-While the generalization error for arbitrary distributions of neuronal activities and latent variables is analytically intractable, we show that in many cases, the error only depends on a few key statistics.
-To do this, we draw from recent work in deep learning theory showing that the generalization error of linear readouts trained on complex distributions can in many cases be well approximated by studying simpler Gaussian models (Methods) [30, 35, 31, 36].
-Thus, we analytically calculate the generalization error using a Gaussian model and show empirically that our theory accurately accurately predicts the generalization error of the linear readout rule when applied to non-Gaussian synthetic data (Sec.
-3.3) as well as real neural responses from V4 and IT on a visual classification task (Sec.
-3.4).
-Using this simplified model, we prove a formula for the generalization error.
-As described in the Supplemental Material (SM), we show that under certain limiting conditions, the generalization error $E_{g}$ can be written as a strictly decreasing function of 3 geometric terms, which we schematize in Fig.
+(b) Visualization of neural activity patterns with low (left) and high (right) total correlation.
+When the correlation is high, the relative distances between points in the latent space are approximately preserved in the neural state space.
+(c) Signal-signal factorization (SSF).
+When the SSF is low, different latent variables are represented along overlapping directions, and when it is high, independent directions in the latent space are represented along approximately orthogonal directions in the neural state space.
+(d) Signal-noise factorization (SNF).
+When the SNF is low, the noise distribution (grey ellipses) around a point in the firing rate space falls along the coding directions.
+When it is high, the noise distribution is orthogonal to these directions.
+(e) Neural dimension.
+In higher dimensional representations, the neural activity and associated noise distribution occupies more directions in the state space, shown here as 2d (left) vs.
+3d (right) noise distributions.
+As the dimension increases, the projection of a sample of neural activity onto a given direction becomes increasingly concentrated, supporting generalization performance [26].
+Using this simplified model, we derive a formula for the average generalization error.
+As described in the Supplemental Material, we prove that under certain conditions, the generalization error can be written as a strictly decreasing function of four geometric terms, which we schematize in Fig.
 2.
@@
   • Neural-latent correlation: $c.$ This term is a normalized sum of all covariances between neurons and latent variables.
-Geometrically, the total correlation measures how sensitive the population responses are to variations in the latent space.
-  • Alignment: $a.$ This term measures the alignment between the coding directions for each latent variable and the total variation in the neural response.
-Because neural response variation reflects a combination of latent-driven variation (i.e., signal) and latent-independent noise, the alignment term favors encoding each latent variable along a direction that contains little signal from the other latent variables and which has low latent-independent noise.
-Thus, a high alignment $a$ reflects low noise in the coding directions and near orthogonal coding directions for each latent variable (i.e., a factorized code).
-  • Neural Dimensionality: $\operatorname{PR}.$ The participation ratio of the neural responses measures the number of dimensions that the population activity spans.
+Geometrically, the neural-latent correlation measures how sensitive the population responses are to variations in the latent space.
+  • Signal-signal factorization (SSF): $f$.
+This term measures the alignment between the coding directions of distinct latent variables.
+The signal-signal factorization term favors neural coding schemes that represent distinct latent variables along orthogonal directions in the state space.
+Moreover, this term encourages these representations to devote equal variance to distinct coding directions–i.e., to form a whitened representation of the latents.
+  • Signal-noise factorization (SNF): $s$.
+This term measures the magnitude of the noise that lies along the coding directions of the latent variables.
+Ideally, any noise present in the neural responses should lie in directions which are orthogonal to the directions representing the latent variables.
+  • Neural dimension: $\operatorname{PR}.$ The participation ratio of the neural responses measures the effective number of dimensions that the population activity spans.
 When all else is equal, higher dimensional responses are preferred, as neuronal noise is less correlated from trial to trial [26].
-Finally, we note that the contribution to the error of neuronal dimensionality $\operatorname{PR}$ decays to 0 with the number of training samples (Methods).
-Similarly, we find that the correlation term’s contribution to the error decays with the number of samples.
-On the other hand, the error stemming from the alignment, $a,$ represents an irreducible error that does not depend on the number of training samples.
-Thus, we anticipate that as the number of training samples grow, it becomes increasingly important to maximize the alignment term, while in the few-shot regime, the correlation and participation ratio terms dominate the error.
+Finally, we note that the contribution to the error of the neural dimension and neural-latent correlation decays to 0 with the number of training samples (Methods).
+On the other hand, the error stemming from the signal-signal and signal-noise factorization represents an irreducible error that does not depend on the number of training samples (SM).
+Intuitively, in the few-shot regime, the main concern is maximizing the amount of total signal in the neural code, while minimizing the impact of noise.
+These two aspects are primarily controlled by the total correlation and dimension, respectively.
+On the other hand, in the many-shot regime the main concern becomes keeping the representations of distinct latent variables separate from one another and separate from noise directions.
+These features of the code are respectively controlled by the two factorization terms.
 Figure 3: Theory predicts empirical generalization error in Gaussian model with power law covariance spectra.
@@
 We set the covariance matrix of the latents to have a spectrum that decays as a power law.
-Since the proof of our main theorem assumes that the spectra of the covariances decays more slowly than a power law with rate 1, this allows us to parametrically study violations of our main theorem’s assumptions (Fig.
-3a&e; SM).
+Since the proof of our main theorem assumes that the spectra of the covariances decays slowly, this allows us to parametrically study violations of our main theorem’s assumptions (Fig.
+3a-e; SM).
 After sampling the latents, we form the neural responses either by taking a random high dimensional projection of the latent variables, or by applying a whitening transform to the latent variables (Fig.
@@
 3f-h, our theory yields an excellent fit to numerical simulations.
-Importantly, the theoretical prediction holds all the way down to the few shot learning regime in which $p$ is small, despite the theory being derived in the limit of large $p$.
-Furthermore, we find that the theory predicts the empirical generalization error well for relatively small values of $n$ and $d$, as well as latent variables whose covariance has an eigenspectrum that decays faster than what our main theorem assumes ($\alpha<1$, where $\alpha$ is the rate of decay).
+Importantly, the theoretical prediction holds all the way down to the few shot learning regime in which the number of training samples $p$ is small, despite the theory being derived in the limit of large $p$.
+Furthermore, we find that the theory predicts the empirical generalization error well for a relatively small number of neurons $n$ and latent dimensionality $d$, as well as latent variables whose covariance has an eigenspectrum that decays relatively quickly [32] (Methods).
 These results suggest that our theory can be applied to a wide range of datasets and is informative of the few shot learning regime.
@@
 (a) In our task setup, directions in the latent space that have little variance are, on average, less informative of the task labels (Methods).
-Optimal neuronal representations of these latent variables map independent directions in the latent space to independent directions in the neuronal space–i.e.
-they are factorized.
+Optimal neuronal representations of these latent variables map independent directions in the latent space to independent directions in the neuronal space.
 The amount of variance corresponding to less informative directions in the latent space is small when data is scarce and is large when data is abundant.
 (b) Eigenvalues of the optimal neural covariance as a function of the number of samples vs.
-latent dimensionality, $p/d$.
+latent dimension, $p/d$.
 We show the eigenvalues of the latent variables’ covariance in black.
@@
 As the number of samples increases, the spectrum of the optimal neural code becomes increasingly flat, indicating the expansion strategy described above.
-(d-f) Geometric terms of the optimal neural representation for various values of $p$ and using the same latent covariance as in (c).
-Surprisingly, the total correlation decreases with $p$.
+(c-e) Geometric terms of the optimal neural representation for various values of $p$, given the same latent covariance.
+(Note that we do not plot the signal-noise factorization, as it diverges for the optimal representation for all $p$.) Surprisingly, the total correlation decreases with $p$.
 Our theory provides an ideal framework to pose normative questions regarding multi-task learning.
@@
 4a).
-This allows us to probe how task structure and the number of available training samples determine the optimal representation.
-In this section, we study what the best possible neural representational geometry is, given some latent variables and a fixed training dataset size.
+This allows us to probe how task structure and the number of available training samples determine which neural representations are optimal in the sense of achieving the lowest possible generalization error.
+In this section, we study the geometry of optimal neural representations, given some latent variables and a fixed training dataset size.
 We show analytically that the optimal representation disentangles latent variables into orthogonal subspaces.
@@
 As shown in Fig.
-4a, when the training budget is small, less informative latent variables correspond to directions in the neuronal activity space that have small variance.
-As $p$ increases, the amount of variance dedicated to these less informative latent variables grows.
+4a, when training data is limited, less informative latent variables correspond to directions in the neuronal activity space that have small variance.
+With increasing training data (i.e., as $p$ increases), the amount of variance dedicated to these less informative latent variables grows.
 This reflects a strategy in which the optimal code compresses less informative variables at small training set sizes $p$, while at large sample sizes, these variables are expanded in the state space (Methods).
 Speaking informally, the optimal code only starts paying attention to the less informative latent variables when there are enough samples to learn their relevance to a given task.
-We trace these features of the optimal code back to the eigenvalues of the neuronal covariance matrix and our three geometric terms.
+We trace these features of the optimal neural code back to the eigenvalues of the neuronal covariance matrix and our geometric terms.
 As shown in Fig 4b, the eigenspectrum of the optimal code becomes increasingly flat as $p$ increases.
 This reflects the fact that more and more variance is being dedicated to the less informative directions in the latent space (i.e., directions with smaller variance).
-Turning to the geometric terms, we can see that tradeoff appears as a trend in which the participation ratio and alignment of the optimal code both increase with $p$, while the correlation decreases (Fig.
+Turning to the geometric terms, we can see that this trade off appears as a trend in which the participation ratio and signal-signal factorization term of the optimal code both increase with $p$, while the correlation decreases (Fig.
 4c-e).
-The fact that high dimensional representations can be preferred for large $p$ is a surprising consequence of our calculation, as the neuronal dimensionality becomes less important to the error as $p$ grows (Methods).
-A key normative prediction from our theory is therefore that the spectrum of neural responses becomes increasingly high dimensional as an agent learns to perform tasks that depend on a complex latent structure.
+That is, optimal neural codes become increasingly high dimensional as the number of available training samples increases.
+A key normative prediction from our theory is therefore that the task-related variability of neural responses becomes increasingly high dimensional as an agent learns to perform tasks that depend on a complex latent structure.
 3.3 Geometry of multi-task learning in MLPs
@@
 Theoretical predictions closely track empirical errors, and the trained network achieves a lower error in intermediate layers.
-(d-f) Geometric terms across layers for the random (light blue line) and trained (dark blue line) networks.
+(d-g) Geometric terms across layers for the random (light blue line) and trained (dark blue line) networks.
 Linear layers are marked by circles and relu layers by squares.
 Interestingly, the error only slightly changes across linear and relu layers of the same model stage, in spite of sharp changes in the geometry.
-(g-i) Average change in the geometry between relu and linear layers.
+(h-k) Average change in the geometry before and after the application of relu.
 Here we show the mean and standard deviation of the difference between each geometric term before and after the relu nonlinearity is applied.
-The application of relu consistently causes increases in the dimension and alignment, as well as decreases in the correlation.
-To test our theory on explicitly non-Gaussian data, we now apply our theory to both random and trained non-linear multi-layer perceptrons (MLPs).
+In the trained network, the application of relu consistently causes increases in the dimension and signal-signal factorization, as well as decreases in the correlation.
+In practice, neural data is complex and may be non-Gaussian.
+To test our theory in this regime, we now apply our theory to both random and trained non-linear multi-layer perceptrons (MLPs).
 As shown in Fig.
@@
 These labels and data points from the random MLP are then used to train a downstream 4-hidden-layer MLP, which predicts the label for each task using a linear readout of its penultimate layer.
-This is reminiscent of the hidden manifold modeling framework studied in the machine learning community [30].
+This is reminiscent of the hidden manifold modeling framework [36].
 Finally, we sample a new set of latent variables along with a new set of tasks and calculate the generalization error of the Hebb rule when applied to the representations at different layers of the trained and untrained MLPs (Methods).
@@
 Turning to our geometric decomposition of the error, we find several interesting trends across layers and non-linearities (Fig.
-5d-f).
-In the random MLP, we see that the correlation decreases throughout the network, while the alignment term exhibits small fluctuations while gradually decreasing overall.
-Since we choose the layer sizes to be monotonically increasing through the network, we also find that the dimension of the responses stays more or less constant through the network.
-In the trained network, we find that linear and relu layers orchestrate a trade off between dimensionality, alignment, and correlation that together only slightly perturb the generalization error.
-As shown in Fig 5g-i, the neural dimensionality and alignment spike each time the nonlinearity is applied, at the cost of the total correlation.
-Conversely, at each linear layer, the correlation sharply rises while the dimensionality and alignment fall.
+5d-g).
+In the random MLP, we see that the correlation decreases each time the non-linearity is applied, while the signal-noise and signal-signal factorization terms gradually decrease through the network.
+We also find that the dimension of the responses stays more or less constant through the network.
+In the trained network, we find that linear and relu layers orchestrate a trade off between the geometric terms that together lead to an overall decrease in multi-task generalization error.
+As shown in Fig 5h-k, the neural dimension spikes each time the nonlinearity is applied, at the cost of the total correlation.
+Conversely, at each linear layer, the correlation sharply rises while the dimension falls.
+On the other hand, we find that both the signal-signal and signal-noise factorization terms monotonically increase through the penultimate layer of the network.
 Importantly, this pattern is not present in the random network.
-Thus, the trained MLP learns to use the nonlinearity to increase the dimension of the representations while simultaneously squashing particularly harmful noisy directions of variability in the input and increasing how well factorized the code is.
+Thus, the trained MLP learns to use the nonlinearity to increase the dimension of the representation as well as the factorization of the latent variables, while simultaneously squashing particularly harmful noisy directions of variability in the input.
 Intuitively, the linear layers learn to orient signal-unrelated features into parts of the state space that are zeroed out by the non-linearity [43].
@@
 6, we find that the generalization error in this context decays through training across layers.
-In the linear layers, we find a small increase in the total correlation and dimension as well as a steady rise in the alignment.
-On the other hand, in relu layers, improved generalization is driven by an increase in the dimension and alignment across both early and late layers.
-In these layers, the total correlation decreases throughout training.
-We find similar trends for MLPs with a $\tanh$ non-linearity (SM).
-These geometric trends mirror the effects of adding more samples on the optimal representation: As more samples become available, the alignment and dimension grow while the correlation shrinks.
+In the linear layers, we find that improved generalization is driven by a steady rise in the signal-signal and signal-noise factorization as well as small increases in the total correlation and dimension.
+On the other hand, in relu layers, the correlation decreases through training, with improved generalization being driven by increases in the dimension as well as the signal-signal and signal-noise factorizations across both early and late layers.
+We find similar trends for MLPs with a $\tanh$ non-linearity (SM Figs.
+2&3).
+These geometric trends mirror the effects of adding more samples on the optimal representation: As more samples become available, the factorization terms and dimension grow while the correlation shrinks.
 In other words, the dynamics of MLP representations in non-linear layers qualitatively mimic the behavior of the optimal neural code described by our theory.
@@
 This term decreases through training for relu layers and slightly increases for linear layers.
-(d) Alignment term.
-This uniformly increases through training across the network.
+(d) Signal-signal factorization term.
+(e) Signal-noise factorization term.
+Both of these terms uniformly increase through training across the network, though the former begins to plateau at the end of training.
 These findings mirror the effect of increasing the number of training samples on the optimal neural code.
@@
 [44].
-Applying our geometric decomposition of the generalization error to these data, we find that the dimensionality is lower in V4 than in either IT or the pixel data [22] (Fig.
+Applying our geometric decomposition of the generalization error to these data, we find that the dimension is lower in V4 than in either IT or the pixel data [26] (Fig.
 7e).
-Furthermore, we find that the alignment is far lower in V4 than in IT, suggesting that factorization of these latent variables increases along the ventral stream, in line with previous results [39].
-However, we find that the alignment is higher in the pixels than in either V4 or IT.
-This finding suggests that the ventral stream representations considered here do not factorize information regarding object position well, as a poor alignment indicates that neuronal noise and the coding directions of distinct latents are not orthogonal to one another (Fig.
-7e-g).
-Nevertheless, we do find that the total correlation is much higher in V4 and IT than in the pixel space.
-This suggests that while the representations considered here are not well factorized, they do contain correlations with the latent variables that can be leveraged to achieve a generalization error that is better than classification on the pixels.
+Furthermore, we find that the correlation steadily increases from the raw pixels to V4 and IT, indicating an increased overall signal.
+We also find that the level of signal-signal factorization improves from V4 to IT, in line with previous results [45].
+Interestingly, we find that the signal-signal factorization is higher in the pixel space than in either brain region.
+Finally, we find that the signal-noise factorization is far lower in the pixels and V4 than in IT.
+This suggests that in IT, latent-unrelated variability overlaps less with the coding directions than in V4 or the pixels (Fig.
+7e-h).
 Taken together, these results provide strong support for our theory and demonstrate the applicability of our metrics as a tool for tying the geometry of neuronal population responses to the computational objective of multi-task learning.
@@
 (d) Generalization error across pixels and brain regions calculated empirically (markers) and using our formula for the generalization error (solid line).
-(e-g) Geometric terms across pixels and neural responses.
-We can see that V4 representations are lower dimensional than either the pixels or IT, and that both IT and V4 suffer from a lower alignment than the pixels.
-However, both regions present with a much higher total correlation with the latent variables, enabling improved generalization performance.
+(e-h) Geometric terms across pixels and neural responses.
+We can see that V4 representations are lower dimensional than either the pixels or IT, that the correlation improves throughout, and that the signal-signal factorization is highest in the pixel space, though it improves from V4 to IT.
+Note too that the signal-noise factorization sharply rises from V4 to IT, suggesting that signal-unrelated variability overlaps less with the coding directions in the latter region.
 4 Discussion
@@
 In this model, binary classification tasks are generated by shattering a latent space uniformly.
-By varying the spectrum of the latent covariance, we were able to model situations in which certain latent variables are, on average, more useful for downstream tasks than others.
+By varying the spectrum of the latent covariance matrix, we were able to model situations in which certain latent variables are, on average, more useful for downstream tasks than others.
 We then calculated the generalization error of a linear readout applied to a set of neural activations responding to stimuli that were generated from these latent variables.
 In this way, we evaluated how well a neural representation is able to extract information about latent variables from the stimuli and use this for arbitrary downstream tasks.
-To connect the mesoscopic statistics of neural activation patterns to generalization performance, we decomposed our formula for the generalization error into three geometric terms.
+To connect the mesoscopic statistics of neural activation patterns to generalization performance, we decomposed our formula for the generalization error into four geometric terms.
 These terms summarize the relationship between population activity and the latent variables and completely determine generalization performance in our task.
-First is the dimensionality of the entire population activity, as measured by the participation ratio.
+First is the dimension of the entire population activity, as measured by the participation ratio [46].
 Second, we introduced the total correlation between neuronal activity and the latent variables, representing the overall sensitivity of neuronal responses to changes in the latent space.
-Finally, we defined the alignment, which summarizes the degree to which distinct latent variables and latent-unrelated variability are coded along orthogonal directions in the neural state space.
-As such, this term measures the degree to which a representation is factorized and manages to keep noise variability out of the coding directions.
-Together, these three terms completely characterize the generalization performance of the linear readout applied to the neural population responses.
-To investigate the role that latent structure and data scarcity play in shaping population codes, we next analytically calculated which geometries minimize the generalization error, given a fixed latent structure and number of available training samples.
+Third, we introduced the signal-signal factorization, measuring the degree to which each latent variable is coded in an orthogonal direction in the neural state space, with adequate variance devoted to each coding direction.
+Finally, we defined the signal-noise factorization, which summarizes the amount of signal-unrelated variability that lies along the directions in state space used to represent the latent variables.
+Hence, this term summarizes the degree to which noise correlations affect the readout considered here [47, 48, 49, 50].
+Together, these four terms completely characterize the generalization performance of the linear readout applied to the neural population responses.
+We next analytically calculated which geometries minimize the generalization error, given a fixed latent structure and number of available training samples.
 This falls in line with studies on efficient coding in which predictions of sensory physiology are derived from statistics of the natural environment [51, 52, 53].
-In this way, we found that disentangled (or factorized) representations naturally emerge as an optimal solution to the multi-task learning problem.
-Specifically, we found that the eigenvectors of the neuronal and latent covariances match the singular vectors of the cross-covariance (Methods).
+In this way, we found that disentangled representations naturally emerge as an optimal solution to the multi-task learning problem.
+Specifically, we found that the eigenvectors of the neuronal and latent covariances match the singular vectors of the cross-covariance (Methods, SM).
 This is reminiscent of work done on the successor representation and grid cells [54, 55, 56, 57, 58], and it would be interesting to see if there are additional constraints under which grid cells or other coding properties emerge as an optimal code in our model [59].
+Future work can also extend our theory to consider sets of tasks that compete with one another as in previous studies of multi-task learning [60, 61].
+Moreover, given recent findings regarding multi-task learning and dynamical motif reuse in recurrent networks [62, 63], it would be interesting to extend our theory to a dynamical setting in which predictions are formed using an appropriate readout of time-varying firing rates.
 We then described how the optimal neuronal code shifts from being low to high dimensional as the number of available samples increases.
 This finding reflects a strategy of compressing less useful information when data is scarce and expanding it when data is abundant.
-It is possible that a similar relation holds for low vs.
-high noise learning scenarios [49, 50, 51, 52].
-It is interesting to compare these findings to previous work [22] reporting that higher dimensional representations of individual object classes are preferred for invariant object recognition in the few shot learning setting.
-Although the contribution to the error of the dimensionality scales as 1/ p in our formula just as in this work, the additional contribution to the error of the alignment term, leads to optimal representations with lower dimensionality in the few shot regime.
+It would be very interesting to test this normative prediction in behaving animals.
+This could be done, for example, using multi-unit recordings taken while an animal learns to perform decision making tasks in which multiple latent variables (e.g.
+distinct context cues) must be used to make a decision.
+Our theory predicts that task-related neural variability in the early stages of training should primarily track highly informative latent variables.
+As training progresses, we predict that task-relevant neural variability will increase in dimension as it begins to track the less informative latent variables in the task.
+It is interesting to compare these findings to previous work reporting that higher dimensional representations of individual object classes are preferred for invariant object recognition in the few shot learning setting [26].
+Although the contribution to the error of the neural dimension scales as 1/ p in our formula just as in this work, the additional contribution to the error of the signal-signal and signal-noise factorization terms together with the constraint of of the covariance matrix remaining positive semi-definite, leads to optimal representations with lower dimension in the few shot regime.
 These results highlight the fact that different geometric terms may compete with each other in ways that cannot be directly read off from generalization error equations when there are additional constraints imposed on the system.
-Finally, we note that theories of least squares estimation suggest that similar phenomena occur for more complex readouts [53, 54, 55, 56].
-We suspect that the same spectral expansion trend we describe holds for such readouts, though perhaps the optimal code can remain smooth in such cases [27].
-It would be interesting to investigate these questions either analytically or numerically in future work.
+Another related line of work has considered the role of neural dimension in random pattern separation [64, 65, 66].
+In these settings, the optimal dimension can decrease when neural noise increases, and future work can examine whether a similar relation holds in the multi-task learning setting considered here.
+Finally, we note that theories of least squares estimation suggest that similar phenomena occur for regression tasks with more complex readouts [67, 68, 69, 70].
+We suspect that a similar spectral expansion trend of optimal coding schemes holds for other common readout mechanisms in both classification and regression tasks, and it would be interesting to investigate these questions in future work.
 To test our theory, we next considered a setting in which a set of Gaussian latents are fed through a random network, and task labels are generated as above in the latent space.
@@
 Applying our geometric decomposition of the error, we found that linear and relu layers obtained similar generalization errors but with sharply different underlying geometries.
-Specifically, relu layers yielded a large boost to dimensionality and alignment at the cost of the total correlation.
+Specifically, relu layers yielded a large boost to the dimension at the cost of the total correlation, while the signal-signal and signal-noise factorizations steadily improved through the network.
 This suggests a strategy in which linear layers attempt to place neuronal noise in regions of the state space that are then squashed by the non-linearity [43].
-Tracking the evolution of these terms through training, we found that the correlation in relu layers decreases through training, while the alignment and dimension increase.
+Tracking the evolution of these terms through training, we found that the correlation in relu layers decreases through training, while the dimension, signal-signal factorization, and signal-noise factorization increase.
 This mirrors the behavior of an optimal code as more samples are presented as well as the behavior of deep linear networks trained on certain regression tasks [71, 72].
@@
 We found that both regions showed better readout performance than the pixels themselves [44].
-While we found a relatively low alignment in these neural responses, we did find that alignment improved from V4 to IT.
-These findings suggest that, while these latent variables are not overall well factorized in the ventral stream, the degree of factorization improves from V4 to IT, partially in line with previous results [39].
-It would be very interesting to repeat this experiment with visual regions specialized for localization.
+While we found a low signal-signal factorization in these neural responses relative to the pixels, we did find that signal-signal factorization improved from V4 to IT [45].
+Moreover, we found a sharp increase in the signal-noise factorization from the pixels and V4 to IT, suggesting that latent-unrelated variability may become increasingly orthogonal to the coding directions through the ventral stream.
+It would be very interesting to repeat such an analysis using data from visual regions specialized for localization.
 Overall, these results demonstrate the applicability of our theory as a data-analytic tool.
-While we primarily focused on neural representations, it would be very interesting to derive behavioral predictions from this theory.
-Specifically, one could decompose the generalization error formula into error rates for easy vs.
-hard examples.
-In this way, we could leverage our theory to predict learning curve dynamics across stimuli of varying levels of difficulty.
-These predictions could then be tested in perceptual discrimination tasks involving continuous modulation of an unobserved latent space.
+While we focused on tasks that came from shattering a continuous latent space, future work could extend this theory to consider distributions of latents and tasks that more closely mirror common experimental settings.
+Specifically, one could repeat our general calculation, while restricting the distribution of task vectors to a handful of relevant directions or restricting the distribution of latent variables to be fixed to discrete values.
+This would be particularly interesting in settings where only particular groups of tasks or latent variable values are relevant for an animal to consider.
+Such a calculation would allow one to tie the population geometry to the linear readout performance on families of tasks and latent distributions which are tailored to specific experimental settings.
 We hope to pursue this line of research in subsequent work.
@@
 Associated with each latent, we also consider $n$-dimensional neural activity patterns, $x_{\mu}\in\mathbb{R}^{n}$.
-From these data, the agent then forms predictions of new data points using a supervised Hebbian readout rule [59, 32].
+From these data, the agent then forms predictions of new data points using a supervised Hebbian readout rule [38].
 More precisely, given a new stimulus associated with latent variables $z_{+}$ and firing rates $x_{+}$, the agent forms a prediction $\hat{y}_{+}$ using the rule
@@
 1f).
-We evaluate how well the neural code can support downstream classification by calculating the generalization error of these predictions across different tasks (i.e., different $T$ vectors).
+We evaluate how well the neural code can support downstream classification by calculating the generalization error of these predictions, averaged across different tasks (i.e., different $T$ vectors), though we also calculate the error for a fixed $T$ along the way (SM).
 While we validate our theory on a wide range of data, we obtain our analytical results using a simplified Gaussian model.
-This is inspired by work on the Gaussian Equivalence Principle (GEP) in the deep learning theory community [35, 30, 31].
+This is inspired by work on the Gaussian Equivalence Principle (GEP) in deep learning theory [40, 36, 37].
 Formally, we assume that each pair of neural responses and latent variables, $(x,z)$ are jointly zero-mean Gaussians with covariance matrices:
@@
 To do so, we consider the limiting case in which the number of neurons, latent dimensions, and training samples are all large and of comparable size.
-Furthermore, we assume that the covariance matrices given above satisfy certain spectral properties–essentially we assume that their spectra decay at least as slowly as a power law with rate $\alpha<1$ (SM).
-Under these assumptions, we show that the average generalization error, $E_{g}$, is a decreasing function of the three geometric terms.
+Furthermore, we assume that the covariance matrices given above satisfy certain spectral properties (SM).
+Under these assumptions, we show that the average generalization error, $E_{g}$, is a decreasing function of the four geometric terms.
 Formally, we have
-$E_{g}=\frac{1}{\pi}\tan^{-1}\bigg{(}\sqrt{\frac{\pi}{2pc^{2}\mathrm{PR}(\Psi)}+\frac{1}{ca}-1}\bigg{)},$ (3)
-where we have introduced the total neural-latent correlation $c$, the alignment $a$, and the dimensionality of the population responses as measured by the participation ratio, $\mathrm{PR}(\Psi)$.
-These three terms correspond to statistics of the covariance matrices specified above:
-$c$ $=\frac{\mathrm{Tr}(\Phi\Phi^{T})}{\mathrm{Tr}(\Psi)\mathrm{Tr}(\Omega)}$ (4)
-$a$ $=\frac{\mathrm{Tr}(\Phi\Phi^{T})\mathrm{Tr}(\Psi)}{\mathrm{Tr}(\Phi^{T}\Psi\Phi)}$ (5)
+$E_{g}=\frac{1}{\pi}\tan^{-1}\bigg{(}\sqrt{\frac{\pi}{2pc^{2}\mathrm{PR}(\Psi)}+\frac{1}{f}+\frac{1}{s}-1}\bigg{)},$ (3)
+where we have introduced the total neural-latent correlation $c$, the signal-signal factorization $f$, the signal-noise factorization $s$, and the dimension of the population responses as measured by the participation ratio, $\mathrm{PR}(\Psi)$.
+These terms correspond to statistics of the covariance matrices specified above:
+$c$ $=\frac{\mathrm{Tr}(\Phi\Phi^{\top})}{\mathrm{Tr}(\Psi)\mathrm{Tr}(\Omega)}$ (4)
 $\operatorname{PR}(\Psi)$ $=\frac{\mathrm{Tr}(\Psi)^{2}}{\mathrm{Tr}(\Psi^{2})}$ (5)
-Since the function $F(w)=\frac{1}{\pi}\tan^{-1}\big{(}\sqrt{x-1}\big{)}$ is strictly increasing, we can see that each of these three geometric terms should be made as large as possible.
+$f$ $=\frac{\mathrm{Tr}(\Phi\Phi^{\top})^{2}}{\mathrm{Tr}(\Omega)\mathrm{Tr}(\Phi^{\top}\Phi\Omega^{-1}\Phi^{\top}\Phi)}$ (6)
+$s$ $=\frac{\mathrm{Tr}(\Phi\Phi^{\top})^{2}}{\mathrm{Tr}(\Omega)\mathrm{Tr}(\Phi^{\top}(\Psi-\Phi\Omega^{-1}\Phi^{\top})\Phi)}$ (7)
+Since the function $F(w)=\frac{1}{\pi}\tan^{-1}\big{(}\sqrt{w-1}\big{)}$ is strictly increasing, we can see that each of these geometric terms should be made as large as possible.
+As discussed in the SM, the term $f$ measures the overall degree of orthogonality between coding directions, while $s$ measures the amount of noise that lies along the signal directions.
+The neural noise is described by the noise covariance matrix, $\mathrm{cov}(x|z)=\Psi-\Phi\Omega^{-1}\Phi^{\top},$ which appears in the definition above.
+Importantly, this neural noise may include variability that is related to latent variables that are not measured experimentally.
+Motivated by this observation, we describe how these two factorization terms can be collapsed into a single factorization term in the SM.
 Note that in writing Eq.
-(3), we have separated a term that depends on the number of training samples, $1/[c^{2}\operatorname{PR}(\Psi)]$ from a term which is independent of the sample size, $1/(ca)$.
-Thus we can see that the correlation and dimensionality terms become less important as $p$ grows.
+(3), we have separated a term that depends on the number of training samples, $1/[c^{2}\operatorname{PR}(\Psi)]$ from a term which is independent of the sample size, $1/f+1/s$.
+Thus we can see that the correlation and dimension terms become less important as $p$ grows.
 5.3 Gaussian simulations
@@
 Using this approach, we find that the left and right singular vectors of $\Phi$ are the eigenvectors of $\Psi$ and $\Omega$, respectively for the optimal code.
-This shows that independent directions in the latent space map directly onto independent directions in the firing rate space.
+This shows that independent directions in the latent space map directly onto independent directions in the firing rate space (SM).
 Furthermore, we obtain the following simple formula for the eigenvalues of the optimal neural code.
 Denoting the eigenvalues of $\Psi$ and $\Omega$ as $\psi_{i}$ and $\omega_{i}$ respectively, we have up to a permutation symmetry:
-$\psi_{i}=C\frac{\omega_{i}}{2p\omega_{i}+\sum_{k}\omega_{k}},$ (7)
+$\psi_{i}=C\frac{\omega_{i}}{2p\omega_{i}+\pi\sum_{k}\omega_{k}},$ (8)
 where $C$ is an arbitrary constant.
@@
 We used random and trained MLPs to test several predictions from our theory using explicitly non-Gaussian artificial neural response data.
-To generate these data, we first sampled a set of $d=40$ dimensional latent variables from a multivariate Gaussian distribution with eigenvalues $\omega_{k}=k^{-0.2}$, though we also show similar results for a flat spectrum (SM).
+To generate these data, we first sampled a set of $d=40$ dimensional latent variables from a multivariate Gaussian distribution with eigenvalues $\omega_{k}=k^{-0.2}$.
 For these experiments, we sampled $5\cdot 10^{5}$ latent variables independently from this distribution as a training dataset.
@@
 We chose the size of each intermediate layer to be twice the size of the previous one to ensure the dimension of the representation did not decrease.
-Each layer was composed of a linear transform, followed by a relu non-linearity (see SM for results using a tanh non-linearity).
+Each layer was composed of a linear transform, followed by a relu non-linearity (see SM Figs.
+2&3 for results using a tanh non-linearity).
 After sampling the task labels and passing the latents through the random MLP, we trained a $4$ hidden-layer MLP to predict task labels from the random MLP responses.
 The trained network was similarly composed of linear-batchnorm-relu blocks, and we used the adam optimizer to train the network [73] through a single epoch.
-This setup corresponds to a multi-task version of the hidden manifold modeling framework studied in the machine learning community [30].
+This setup corresponds to a multi-task version of the hidden manifold modeling framework studied in deep learning theory [36].
 We evaluated the generalization error through layers and training using a new set of latents and tasks.
@@
 We drew from a publicly available dataset containing multi-unit recordings from V4 and IT taken from 2 monkeys [39].
-These recordings were taken as the monkeys viewed visual stimuli as described in [33] and contained 88 V4 and 168 IT neural sites.
+These recordings were taken as the monkeys viewed visual stimuli as described in [39] and contained 88 V4 and 168 IT neural sites, though we reproduce all results projecting IT and pixel responses down to 88 randomly chosen dimensions (SM Fig.
+4).
 The stimuli for this task were drawn from a generative model in which a total of 64 objects coming from 8 categories were displayed against varying backgrounds.
@@
 Error bars denote the standard error of the mean for the estimated generalization error across the $c=8$ category types.
-In the SM, we also show the geometry and generalization errors for each of the categories individually, as well as the error obtained by pooling all categories together.
+In SM, we also show the geometry and generalization errors for each of the categories individually (SM Figs.
+5&6), as well as the error obtained by pooling all categories together (SM Fig.
+7).
 Acknowledgements
-We thank Thomas Yerxa for comments on this manuscript.
-This work was funded by the Center for Computational Neuroscience at the Flatiron Institute of the Simons Foundation and the Klingenstein-Simons Award to S.C.
-All experiments were performed on the Flatiron Institute high-performance computing cluster.
+We thank Thomas Yerxa, Chi-Ning Chou, Arnav Raha, and Jenelle Feather for helpful discussions and comments on an earlier version of this manuscript.
+This work was funded by the Center for Computational Neuroscience at the Flatiron Institute of the Simons Foundation.
+S.C.
+is also partially supported by a Sloan Research Fellowship, and a Klingenstein-Simons Award.
+All experiments were performed on the high-performance computing cluster at the Flatiron Institute.